OpenStudy (anonymous):

Solve the following equation and show your work. 2x + 2(x + 3) = 2(x + 7)

OpenStudy (anonymous):

4x+6=2x+14 2x=8 x=4

OpenStudy (anonymous):

2x+2(x+3)=2(x+7) 2x+2x+6=2x+14 4x+6=2x+14 4x-2x=14-6 2x=8 x=4